OpenStudy (anonymous):

Escape velocity from sun. Rs=7.0 x 10^5 Ms=2.0 x 10^30 kg. I almost solved problem just cant get the right answer with calculating value please help easy medal (;

OpenStudy (anonymous):

1/2mv^2 - mGMs/Rs= 0

OpenStudy (anonymous):

1/2mv^2=mGMs/Rs

OpenStudy (anonymous):

v^2=2GMs/Rs

OpenStudy (anonymous):

Vesc= [2GMs/Rs]^1/2

OpenStudy (anonymous):

Vesc= {2(2.0 x 10^30)(6.67 x 10^-11)/(7.0 x 10^5)}^1/2

OpenStudy (anonymous):

I keep getting a different value than the books answer );

OpenStudy (anonymous):

When you say Rs does that mean the radius from the center of the sun? and Ms is the mass? I don't understand your variables

OpenStudy (anonymous):

19522880.35 and not sure about untis

OpenStudy (anonymous):

Sorry its an escape of the surface of the sun with given radius of sun and mass of sun. They are rounded but thats the values given in problem.

OpenStudy (anonymous):

Okay so those are the only two variables given right?

OpenStudy (anonymous):

yah

OpenStudy (anonymous):

what units are they given in? meters miles? kg pounds?

OpenStudy (anonymous):

Rs= 7 x 10^5 km Ms= 2.0 x 10^30 kg

OpenStudy (anonymous):

And you're trying to find at what velocity does mass m (our object) have to go to escape the gravitational force of the sun?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what did you put as the gravitational constant of the sun?

OpenStudy (anonymous):

6.67x10^-11

OpenStudy (anonymous):

I think my error may be that the radius needs to be in meters?

OpenStudy (anonymous):

Most likely. You need to change all the variables to similar units. Try that first, and if you still get the wrong answer, just let me know.

OpenStudy (anonymous):

got it